OpenStudy (anonymous):

if f(x) = 2x +3 and g(x) = xsquared then g[f(-3)]=?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

thanks! how?

OpenStudy (anonymous):

\[ g=x^{2}, g[f]=(2x+3)^{2}, g[f(-3)]=(2(-3)+3)^{2}=(-6+3)^{2}=(-3)^{2}=9\]

OpenStudy (anonymous):

Wow thank you sooooo much!